Previous

Calendar

March

2003

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

23

JD54

24

JD55

25

JD56

26

JD57

27

JD58
Site Visit (AT, SpB, GM): Details

Data System problem fixed by Gordon (covar data will be ok after this date).

NCTC logger having problems

some spikes in the LI-6262 data (p.ec goes negative)

disconnected the LI-6262 control cable from the MET logger

28

JD59

No LI-6262 calibrations (no calibrations until 7 Mar)

1

JD60

2

JD61

3

JD62

4

JD63

5

JD64

6

JD65

7

JD66
Site Visit (AT, SpB): Details

Removed 2m T/RH sensor (HMP35D) and brought back to trailer for repair

re-connected the LI-6262 control cable to the MET logger (cals now ok)

8

JD67

9

JD68

10

JD69

11

JD70

12

JD71
Site Visit (SpB): Details

Installed a new 2m T/RH sensor (HMP45D)

Powered off pump for li-6262 (prob. shouldn't have done this!)

13

JD72

14

JD73
Site Visit (SpB): Details

Powered on pump for li-6262

15

JD74

16

JD75

17

JD76

18

JD77
Major Snowstorm (data system down)

19

JD78
Major Snowstorm (data system down)

20

JD79
Data System Down

21

JD80
Spring Equinox
Data System Down

22

JD81
Data System Down

23

JD82
Data System Down

24

JD83

25

JD84
Site Visit (T-seed, SpB, JC): Details

Reboot russter (we could ping it from campus, but could not log in after the snowstorm and power outages)

Replaced GAST pump and NuPro filters in li-6262 system (didn't help increase flow-rate)

26

JD85

27

JD86

28

JD87

29

JD88